OpenStudy (anonymous):

What is the next term of the following sequence? 20, 10, 5, 2 1/2, ...

OpenStudy (anonymous):

1 1/4

OpenStudy (anonymous):

you have to divide by 2 from previous term :)

OpenStudy (anonymous):

I was about to ask that lol

OpenStudy (anonymous):

20/2 = 10 10/2 = 5 5/2 = 1 1/2 and same goes to later :) well, if u're about to ask this, u must be understand d :D

OpenStudy (anonymous):

o opkay